What should be the value of a if the value of 2x 2 + x – a equals to 5, when x = 0?

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

From the question it is given that x = 0
We have,
2x^2+ x–a = 5
a = 2x^2+ x–5
Then, substitute the value of x in the equation
a = (2 × 02) + 0–5
a = 0 + 0–5
a = -5

Thanks